# Primordial black hole formation in $k$-inflation models

### v2

@@ 1 Introduction
An overdense region collapses and forms a PBH when the local value of the density contrast $\delta=\delta\rho/\rho$ is above the so-called collapse threshold density contrast $\delta_{c}$. With the help of Newton’s gravity and Jeans’ radius $R_{\rm J}\sim c_{\rm s}/H$, it was established that PBH forms when the density contrast at the entry of a certain comoving scale $q$ into the acoustic horizon (i.e., when $c_{\rm s}q=aH$) exceeds a threshold value $\delta_{\rm c}\sim c_{\rm s}^{2}$ [7]. Here, $c_{\rm s}$ is the speed of sound equal to $1/\sqrt{3}$ in the case of radiation dominance. The mass of the resulting PBH is approximately equal to the mass within the horizon at the time of formation. Numerical checks for different primordial perturbation profiles show that during radiation dominance, the value for the threshold density contrast $\delta_{c}$ is in the range from 0.40.40.4 to 2/32/32 / 3 [8, 9, 10].
@@
We use the metric signature (+,,,)(+,-,-,-)( + , - , - , - ). Unless specified otherwise, we use the system of units in which $c=\hbar=1$. For convenience, we introduce a length scale $\ell$ so that $\ell\gg 1/M_{\rm Pl}$, where $M_{\rm Pl}=\sqrt{1/(8\pi G)}$ is the reduced Planck mass. By $\mathcal{L}$ and $\mathcal{H}$, we denote the Lagrangian and Hamiltonian multiplied by $\ell^{4}$ so that our $\mathcal{L}$ and $\mathcal{H}$ are dimensionless. By $\varphi$ and $\phi$, we denote the scalar fields connected by $\varphi=\ell^{2}\phi$ and have the dimension of length and mass, respectively.
@@ 2 The PLLS Model
where $\epsilon$ is +1+1+ 1 or 1-1- 1 according to whether $\varphi_{,0}$ is positive or negative. In this model, the variable $X$ can be expressed as an explicit function of $\eta$. By making use of (3) and (10) we find
@@ 4 Curvature perturbations
In the slow-roll regime, the perturbations are conserved once they cross the horizon from the subhorizon to the superhorizon region. A rough estimate [49] shows that the horizon crossing happens at a relatively large $q$-dependent $N$ of the order $N_{q}\sim 7+\ln q/q_{\rm CMB}$, where $q_{\rm CMB}$ is the CMB pivot scale. The curvature perturbations $\zeta_{q}$ are approximately constant at large $N$ once they cross the horizon and enter the superhorizon region where $a(N)H(N)>c_{\rm s}(N)q$, $N>N_{q}$.
@@
[…]
@@
with $-qc_{\rm s}\tau\gg 1$ in the deep subhorizon region. Hence, the initial values of $v_{q}$ and $v_{q}^{\prime}$ at $N_{{\rm in},q}$ are determined by (49) up to an arbitrary phase. The simplest choice is
@@ 4.2 Approximate spectrum
We plot the calculated spectrum for the PLLS model in Fig. 3. In Fig. 4, we plot the spectra calculated for the Tachyon model for two sets of parameters. The approximation (56) correctly reproduces the shape of the spectrum but underestimates the peak by a factor of the order of 101010 and violates the normalization condition (41) (see the note after Eq. (57)).
@@ 5 PBH formation
This power spectrum, which was studied previously in the literature as a toy model (see, e.g., [55]), can serve as a rough approximation to the Tachyon model power spectrum of Fig. 4 with $\mathcal{P}_{0}\simeq 0.1$ 0.20.20.2, $q_{\rm min}=10^{4}q_{\rm CMB}$, and $q_{\rm max}=10^{16}q_{\rm CMB}$.
@@ 5.1 PBH mass fraction
[…]
@@
[…]
@@
[…]
@@
where the height $\beta_{0}$ can be estimated using either (82) or (89). By way of example, we will use $\gamma=0.36$, $\delta_{c}=0.56$, and $\sigma=(4/9)\sqrt{\mathcal{P}_{0}/2}$ with $\mathcal{P}_{0}=0.2$ or 0.40.40.4. Then, using (89) we obtain $\beta_{0}=1.536\cdot 10^{-7}$ for $\mathcal{P}_{0}=0.2$ and $3.429\cdot 10^{-5}$ for $\mathcal{P}_{0}=0.4$. In contrast, using (82) we obtain $\beta_{0}=3.441\cdot 10^{-5}$ and $2.489\cdot 10^{-3}$ for $\mathcal{P}_{0}=0.2$ and $\mathcal{P}_{0}=0.4$, respectively. Hence, the Press-Schechter prediction based on Eq. (82) is by a factor of 100 larger than the prediction of the CCP formalism based on Eq. (89). In the following, we will use the more conservative CCP formalism.
@@ 5.2 PBH abundance
7.5490 0.835370 $-1.4500\times 10^{-8}$ 5.2200 full
PLLS $10^{-16}$ 7.5400 0.835370 $-2.8174\times 10^{-7}$ 5.2500 0.51 dashed
$\alpha=1.5$ 7.5400 0.835370 $-2.8174\times 10^{-7}$ 5.3000 dash-dotted
3.9523 0.895870 $-1.0089\times 10^{-3}$ 5.1124 long dashed
@@
Tachyon $10^{-16}$ 7.5020 0.259500 $-6.1200\times 10^{4}$ 1.1005 0.56 dashed
@@
7.6000 0.259831 $-5.3625\times 10^{4}$ 1.0520 long dashed
@@ 6 Summary and conclusions
[…]
@@ Appendix A The critical collapse threshold
[…]
@@
[…]
@@
[…]

### v3

@@ 1 Introduction
An overdense region collapses and forms a PBH when the local value of the density contrast $\delta=\delta\rho/\rho$ is above the so-called collapse threshold density contrast $\delta_{c}$. With the help of Newton’s gravity and Jeans’ radius $R_{\rm J}\sim c_{\rm s}/H$, it was established that PBH forms when the density contrast at the entry of a certain comoving scale $q$ into the acoustic horizon (i.e., when $c_{\rm s}q=aH$) exceeds a threshold value $\delta_{\rm c}\sim c_{\rm s}^{2}$ [7]. Here, $c_{\rm s}$ is the speed of sound equal to $1/\sqrt{3}$ in the case of radiation dominance. The mass of the resulting PBH is approximately equal to the mass within the horizon at the time of formation. Numerical checks for different primordial perturbation profiles show that during radiation dominance, the value for the threshold density contrast $\delta_{c}$ is in the range from $0.4$ to $2/3$ [8, 9, 10].
@@
We use the metric signature (+,,,)(+,-,-,-). Unless specified otherwise, we use the system of units in which $c=\hbar=1$. For convenience, we introduce a length scale $\ell$ so that $\ell\gg 1/M_{\rm Pl}$, where $M_{\rm Pl}=\sqrt{1/(8\pi G)}$ is the reduced Planck mass. By $\mathcal{L}$ and $\mathcal{H}$, we denote the Lagrangian and Hamiltonian multiplied by $\ell^{4}$ so that our $\mathcal{L}$ and $\mathcal{H}$ are dimensionless. By $\varphi$ and $\phi$, we denote the scalar fields connected by $\varphi=\ell^{2}\phi$ and have the dimension of length and mass, respectively.
@@ 2 The PLLS Model
where $\epsilon$ is +1+1 or 1-1 according to whether $\varphi_{,0}$ is positive or negative. In this model, the variable $X$ can be expressed as an explicit function of $\eta$. By making use of (3) and (10) we find
@@ 4 Curvature perturbations
In the slow-roll regime, the perturbations are conserved once they cross the horizon from the subhorizon to the superhorizon region. A rough estimate [48] shows that the horizon crossing happens at a relatively large $q$-dependent NN of the order $N_{q}\sim 7+\ln q/q_{\rm CMB}$, where $q_{\rm CMB}$ is the CMB pivot scale. The curvature perturbations $\zeta_{q}$ are approximately constant at large $N$ once they cross the horizon and enter the superhorizon region where $a(N)H(N)>c_{\rm s}(N)q$, $N>N_{q}$.
@@
where the conformal time $\tau=\int dt/a$ in the slow-roll regime satisfies
@@
with $qc_{\rm s}\tau\gg 1$ in the deep subhorizon region. Hence, the initial values of $v_{q}$ and $v_{q}^{\prime}$ at $N_{{\rm in},q}$ are determined by (49) up to an arbitrary phase. The simplest choice is
@@ 4.2 Approximate spectrum
We plot the calculated spectrum for the PLLS model in Fig. 3. In Fig. 4, we plot the spectra calculated for the Tachyon model for two sets of parameters. The approximation (56) correctly reproduces the shape of the spectrum but underestimates the peak by a factor of the order of $10$ and violates the normalization condition (41) (see the note after Eq. (57)).
@@ 5 PBH formation
This power spectrum, which was studied previously in the literature as a toy model (see, e.g., [54]), can serve as a rough approximation to the Tachyon model power spectrum of Fig. 4 with $\mathcal{P}_{0}\simeq 0.1$ - $0.2$, $q_{\rm min}=10^{4}q_{\rm CMB}$, and $q_{\rm max}=10^{16}q_{\rm CMB}$.
@@ 5.1 PBH mass fraction
$\beta(M)=\frac{1}{2}{\rm erfc}\left(\frac{\delta_{c}}{\sqrt{2}\sigma}\right)\equiv\frac{1}{\sqrt{2\pi}\sigma}\int_{\delta_{c}}^{\infty}d\delta\exp\left(-\frac{\delta^{2}}{2\sigma^{2}}\right).$ (81)
@@
$\beta(M)=\int_{\delta_{-}}^{4/3}d\delta\frac{M_{\rm PBH}(\delta)}{M}N(\delta),$ (85)
@@
$\beta(M)=\frac{1}{\pi^{2}}\frac{\mu^{3}}{\sigma^{7}}\int_{\delta_{-}}^{4/3}d\delta\left(\delta-\frac{3}{8}\delta^{2}-\delta_{c}\right)^{\gamma}\delta^{3}\exp\left(-\frac{\delta^{2}}{2\sigma^{2}}\right).$ (89)
@@
where the height $\beta_{0}$ can be estimated using either (82) or (89). By way of example, we will use $\gamma=0.36$, $\delta_{c}=0.56$, and $\sigma=(4/9)\sqrt{\mathcal{P}_{0}/2}$ with $\mathcal{P}_{0}=0.2$ or $0.4$. Then, using (89) we obtain $\beta_{0}=1.536\cdot 10^{-7}$ for $\mathcal{P}_{0}=0.2$ and $3.429\cdot 10^{-5}$ for $\mathcal{P}_{0}=0.4$. In contrast, using (82) we obtain $\beta_{0}=3.441\cdot 10^{-5}$ and $2.489\cdot 10^{-3}$ for $\mathcal{P}_{0}=0.2$ and $\mathcal{P}_{0}=0.4$, respectively. Hence, the Press-Schechter prediction based on Eq. (82) is by a factor of 100 larger than the prediction of the CCP formalism based on Eq. (89). In the following, we will use the more conservative CCP formalism.
@@ 5.2 PBH abundance
7.5490 0.835370 $1.4500\times 10^{-8}$ 5.2200 full
PLLS $10^{-16}$ 7.5400 0.835370 $2.8174\times 10^{-7}$ 5.2500 0.51 dashed
$\alpha=1.5$ 7.5400 0.835370 $2.8174\times 10^{-7}$ 5.3000 dash-dotted
3.9523 0.895870 $1.0089\times 10^{-3}$ 5.1124 long dashed
@@
Tachyon $10^{-16}$ 7.5020 0.259500 $6.1200\times 10^{4}$ 1.1005 0.56 dashed
@@
7.6000 0.259831 $5.3625\times 10^{4}$ 1.0520 long dashed
@@ 6 Summary and conclusions
Our results for the PLLS model show that a suitable choice of input parameters different from those in Ref. [20] could yield a significant PBH abundance. For a chosen set of parameters yielding the power spectrum of Fig. 3, the PBH fraction has a few sharp peaks around the PBH mass $M=10^{-11}M_{\odot}$. In contrast, the Tachyon model power spectrum of Fig.  4 predicts experimentally allowed PBH abundance today in a broad mass range (Fig. 6). In the PLLS model, the integrated PBH abundance of the highest peak (full blue line) in Fig. 6 could account for total DM. In the Tachyon model, the highest integrated contribution of PBHs to DM estimate is about $11\%$ (full black line), except for a particular parameterization (long dashed black line) where the integrated PBH contribution may reach up to $98\%$. However, the latter result should not be taken seriously since the spectrum corresponding to this parameterization is not supported by the observational data near the CMB pivot scale.
@@ Appendix A The critical collapse threshold
The conformal time interval $\tau$ is, according to (50), approximately $\tau=-1/aH$. Now we assume that the scale $\hat{r}_{m}$ is a factor of 10 to 100 larger than the comoving horizon radius $1/(aH)$. Hence, in Eq. (98) we may substitute $\hat{r}_{m}/p$ for $\tau$, where $p$ is a constant of order $10\leq p\leq 100$ and, according to [63], the scale $\hat{r}_{m}$ is a positive root of the function
@@
$\alpha_{\rm G}=-\frac{1}{4}\left[1+\dfrac{\int ydy\cos y\,\mathcal{P}_{\rm S}(y/\hat{r}_{m})T^{2}(y/p)}{\int dy\sin y\,\mathcal{P}_{\rm S}(y/\hat{r}_{m})T^{2}(y/p)}\right].$ (100)
@@
$f(\alpha)=\Gamma(5/(2\alpha))-\Gamma(5/(2\alpha),1/\alpha)\equiv\int_{0}^{1/\alpha}dte^{-t}t^{5/(2\alpha)-1}.$ (103)